# Persistent Homology for Structural Characterization in Disordered Systems

### v8

@@ II.1.1 Vietoris-Rips (VR) Complex
The simplices in a complex are not isolated but rather interconnected by sharing vertices and edges, forming a higher-dimensional structure. For instance, If two triangles $[p_{i},p_{j},p_{k}]$ and $[p_{j},p_{k},p_{l}]$ share an edge $[p_{j},p_{k}]$, they are connected to each other by this shared boundary, forming part of the complex. This process generalizes to higher dimensions, yielding a simplicial complex that includes points, edges, triangles, and higher-order simplices. The construction is controlled by $\epsilon$, and different $\epsilon$ corresponds to different geometric realizations of the point cloud, enabling the extraction of its geometric and topological information at varying scales.
@@ II.1.2 Homology Groups
The homology group is a fundamental tool from algebraic topology [54] used to classify topological spaces [66, 67, 68] by analyzing the structure of simplicial complexes. The boundary of a simplex is the most significant to understand the homology. Specifically, a $k$-simplex is a $k$-dimensional analogy of a triangle. For instance, a point is a 0-simplex, an edge is a 1-simplex, a triangle is a 2-simplex, a tetrahedron is a 3-simplex, etc. The boundary of a $k$-simplex is composed of its $(k-1)$-dimensional faces. Each $(k-1)$-dimensional face is defined by a subset of the vertices of the $k$-simplex. For example, the boundary of a solid triangle is all three edges of this triangle, and the face of this solid triangle means the hollow triangle with its interior removed.
The boundary operator $\partial_{k}$ maps each $k$-simplex to its boundary, consisting of $(k-1)$-simplices: $\partial_{k}:C_{k}\to C_{k-1}$. Here, $C_{k}$ represents the $k$-chain group of a chain complex, composed of formal linear combinations of $k$-simplices over a chosen coefficient group, i.e., the cyclic group $\mathbb{Z}_{\mathfrak{p}}$. Specifically, an element $\sigma\in C_{k}$ can be written as a finite sum: $\sigma=c_{1}\sigma_{1}+c_{2}\sigma_{2}+\dots+c_{n}\sigma_{n}$, where $\sigma_{i}$ are $k$-simplices, and $c_{i}\in\mathbb{Z}_{\mathfrak{p}}$ are coefficients subject to modular arithmetic, meaning all addition and scalar multiplication are performed modulo $\mathfrak{p}$. A key property of the boundary operator is that the boundary of a boundary is always zero, i.e., $\partial_{k-1}(\partial_{k}(\sigma))=0$ for any $k$-simplex $\sigma$. This ensures that the boundary of a closed structure, such as a cycle, has no further boundary, which is essential for identifying topological features like cycles.
@@ II.1.3 Persistent Homology
Topological features are “born” when they firstly appear in the complex. For example, a new connected component ($H_{0}$) is born when points first merge as $\epsilon$ increases. Similarly, a cycle ($H_{1}$) is born when a closed cycle (or loop) forms, and a cavity ($H_{2}$) is born when an enclosed space emerges.
@@
In this article, we used the Ripser [75] package to compute persistent homology (PH), and its time complexity for obtaining barcodes from a point cloud with $N$ points, including the computation of $H_{0}$, $H_{1}$, and $H_{2}$ homology classes, is $O(N^{4})$ in the worst case [75].
@@ II.2.3 Identifying Rearrangements
To characterize the tendency of particle rearrangement, we define the positional change $p_{\text{hop},p}(t)$ of a given particle $p$ within a time window $[t-t_{R/2},t+t_{R/2}]$ [79, 80]. If $p_{\text{hop},p}(t)$ exceeds a threshold $p_{c}$ during this interval, the particle is labeled as “soft”; otherwise, it is labeled as “hard” [81]. Specifically, $p_{\text{hop},p}(t)$ is defined by:
@@
where $\mathbb{r}_{p}$ is the position vector of particle $p$, and .$\langle.\rangle_{A}$ and .$\langle.\rangle_{B}$ represent averages over the intervals $A=[t-t_{R/2},t]$ and $B=[t,t+t_{R/2}]$, respectively.
@@ II.3.3 ML Models
    $\mathop{\min}_{\mathbb{w},b}\frac{1}{2}||\mathbb{w}||_{2}^{2}+C\sum_{i=1}^{N}\mathop{\max}(0,1-y_{i}(\mathbb{X}_{(i)}\cdot\mathbb{w}+u))$
@@
    $K(\mathbb{X}_{(i)},\mathbb{X}_{(j)})=\mathop{\exp}(-\gamma||\mathbb{X}_{(i)}-\mathbb{X}_{(j)}||^{2})$
@@ II.3.4 Scalar Field Trained by Linear SVM
$\mathcal{s}_{(p)}(t)=\frac{f(\mathbb{X}_{(p)}(t))}{||\mathbb{w}||_{2}}=\frac{\mathbb{X}_{(p)}(t)\cdot\mathbb{w}+u}{||\mathbb{w}||_{2}}$ (7)
@@
$\mathcal{S}(t)=\frac{1}{\mathscr{N}}\sum_{p=1}^{\mathscr{N}}\frac{\mathbb{X}_{(p)}(t)\cdot\mathbb{w}+u}{||\mathbb{w}||_{2}}$ (8)
@@ II.4 Descriptive Statistics on PH (non-ML approach)
We define a mapping from the results of PH on point cloud data to a real number in the range $[0,+\infty)$, referred to as the Separation Index (SI).
@@
Figure 4: a) is the persistence diagram (PD) for a crystal (Traj($\text{Traj}(T^{(1)}_{5})$ at time step $t=900$), of which the homology classes of $H_{0}$ are distributed along a line near the y-axis, while the $H_{1}$ and $H_{2}$ classes are located within the yellow circle and magnified in b). The yellow solid line in b) divides the point sets of $H_{1}$ and $H_{2}$ into two disjoint parts. Inspired by this, the Separation Index (SI) is defined to measure the clarity of the boundary of the point sets for $H_{1}$ and $H_{2}$.
@@ IV.2 Descriptive Statistics on PH (non-ML approach)
In Sec. II.4, we introduced a mapping from the results of the PH analysis to a single real number in the range $[0,+\infty)$, specifically the Separation Index (SI), as defined in Eq.9. Since both the neighborhood of a particle and the entire system can be represented as point clouds, the SI applies to both local and global structures, mapping each to a single real number.
@@
Figure 7: The $H_{1}$ and $H_{2}$ points for the entire systems from selected trajectories as follow: a) for liquid (Traj($\text{Traj}(T^{(2)}_{9})$ at time step $t=900$); b) for crystal (Traj($\text{Traj}(T^{(1)}_{5})$ at time step $t=900$); c) for amorphous (Traj($\text{Traj}(T^{(2)}_{1})$ at time step $t=200$).
@@ IV.3 Metrics on Trajectories
Figure 9: The a) Separation Index $\text{SI}(t)$ and b) Global Softness $\mathcal{S}(t)$, where $t=100+10\times(i-1)$, $i\in[1,81]\cap\mathbb{N}$, for trajectories in Group 2, denoted as Traj($\text{Traj}(T^{(2)}_{k}),k\in[1,10]\cap\mathbb{N}$. These 10 trajectories are divided into two categories: those with $k\leq 6$ (Category I) crystallized, while those with $k\geq 7$ (Category II) remained liquid. The metrics over time for each trajectory are shown as faded dashed lines, with bold lines indicating the average in each category—blue for Category I and orange for Category II.
@@
We tracked the evolution in four metrics including normalized MSD, $Q_{6}(t)/\mathscr{N}$, $\text{SI}(t)$ and $\mathcal{S}(t)$ over time on ten trajectories in test group (Group 2), i.e., Traj($\text{Traj}(T^{(2)}_{k}),k\in[1,10]\cap\mathbb{N}$, as well as two trajectories in Group 3, i.e., $\text{Traj}_{\text{KA}}^{(1)}$ and $\text{Traj}_{\text{KA}}^{(2)}$. The data for all metrics are provided in Fig. 19 of Appendix E for reference.
@@ A.2 The Parameter Selection for Rearrangement Tendency (Task 1)
Figure 14: The evolution of $p_{\text{hop}}$ for a selected particle (No.310) over a period of time under the lowest temperature (i.e., $T^{(1)}_{1}$). The yellow region represents the interval with the time length of $2t_{R/2}$, encompassing a complete rearrangement, which means $p_{hop}$ rises above $p_{c}$ and then falls back below $p_{c}$. The red-dotted line is the selected $p_{c}=0.1$.
@@
Figure 16: The histogram of the value of $p_{\text{hop}}$ in Traj($\text{Traj}(T^{(1)}_{6})$ and Traj($\text{Traj}(T^{(1)}_{7})$. The onset temperature $T^{(1)}_{c}$ of crystallization separates $T^{(1)}_{6}$ and $T^{(1)}_{7}$, i.e., $T^{(1)}_{6}<T^{(1)}_{c}<T^{(1)}_{7}$. The red-dotted line is the selected $p_{c}=0.1$. The reason for setting $p_{c}=0.1$ is when $p_{\text{hop}}<p_{c}$, the frequency of $p_{\text{hop}}$ in Traj($\text{Traj}(T^{(1)}_{6})$ is larger than that in Traj($\text{Traj}(T^{(1)}_{7})$; when $p_{\text{hop}}\geq p_{c}$, the frequency of $p_{\text{hop}}$ in Traj($\text{Traj}(T^{(1)}_{6})$ is smaller than that in Traj($\text{Traj}(T^{(1)}_{7})$. $p_{c}=0.1$ is the value at which the relative size of two frequency of Traj($\text{Traj}(T^{(1)}_{6})$ and Traj($\text{Traj}(T^{(1)}_{7})$ reverse.
To select $t_{R/2}$, we use the lowest temperature $T^{(1)}_{1}$ as a reference because it corresponds to the system’s lowest mobility. The interval $[t-t_{R/2},t+t_{R/2}]$ should encompass as many complete rearrangements as possible at this low mobility, as shown in Fig. 14. The yellow region marks a full rearrangement, where $p_{\text{hop}}$ rises above and then falls below $p_{c}$. As seen in Fig. 15, most rearrangements at this reference temperature occur within $t_{R}=40$, so we set $t_{R/2}=20$.
As shown in Fig. 16, we chose $p_{c}=0.1$ because it effectively separates the majority of particles in two trajectories above and below the crystallization onset temperature, i.e., $T^{(1)}_{6}$ and $T^{(1)}_{7}$. The $p_{\text{hop},i}(t)$ values are mostly significantly greater or less than $p_{c}=0.1$, respectively. Specifically, $T^{(1)}_{6}$ is the highest temperature at which crystallization occurs, while $T^{(1)}_{7}$ is the lowest temperature where the liquid phase is maintained, both of which are critical for this analysis. By plotting the histograms of $p_{\text{hop}}$ in Traj($\text{Traj}(T^{(1)}_{6})$ and Traj($\text{Traj}(T^{(1)}_{7})$, we determined that $p_{c}=0.1$ is the value at which the relative sizes of the two distributions reverse.
@@ Appendix C Details of Feature Importance Analysis (FIA)
The single decisive predictor mentioned in Sec. IV.1.3, corresponding to the largest eigenvalue from the PCA, is the 1389th element of $\mathbb{I}_{\text{entire}}^{(P)}$, located at pixel (28, 34) in the PI. Given the PI dimensions of $40\times 40$ and birth and persistence ranges of $[-0.1,1.5]$, each pixel has a resolution of $0.04$. We can get that the pixel $(28,34)$ is with birth range of $[1.02,1.06]$ and persistence ranges of and $[0.1,0.14]$ in the PD, respectively. The feature importance, based on the eigenvalues of the principal components, revealed that this key predictor corresponds to the 1389th column of the original feature matrix, which is the 1389th element of the vector $\mathbb{I}_{\text{entire}}^{(P)}$.
@@
[…]

### v9

@@ II.1.1 Vietoris-Rips (VR) Complex
The simplices in a complex are not isolated but rather interconnected by sharing vertices and edges, forming a higher-dimensional structure. For instance, If two triangles $p_{i},p_{j},p_{k}]$ and $p_{j},p_{k},p_{l}]$ share an edge $p_{j},p_{k}]$, they are connected to each other by this shared boundary, forming part of the complex. This process generalizes to higher dimensions, yielding a simplicial complex that includes points, edges, triangles, and higher-order simplices. The construction is controlled by $\epsilon$, and different $\epsilon$ corresponds to different geometric realizations of the point cloud, enabling the extraction of its geometric and topological information at varying scales.
@@ II.1.2 Homology Groups
The homology group is a fundamental tool from algebraic topology [54] used to classify topological spaces [66, 67, 68] by analyzing the structure of simplicial complexes. The boundary of a simplex is the most significant to understand the homology. Specifically, a $k$-simplex is a $k$-dimensional analogy of a triangle. For instance, a point is a 0-simplex, an edge is a 1-simplex, a triangle is a 2-simplex, a tetrahedron is a 3-simplex, etc. The boundary of a $k$-simplex is composed of its (k1)(k-1)-dimensional faces. Each (k1)(k-1)-dimensional face is defined by a subset of the vertices of the $k$-simplex. For example, the boundary of a solid triangle is all three edges of this triangle, and the face of this solid triangle means the hollow triangle with its interior removed.
The boundary operator $\partial_{k}$ maps each $k$-simplex to its boundary, consisting of (k1)(k-1)-simplices: $\partial_{k}:C_{k}\to C_{k-1}$. Here, $C_{k}$ represents the $k$-chain group of a chain complex, composed of formal linear combinations of $k$-simplices over a chosen coefficient group, i.e., the cyclic group $\mathbb{Z}_{\mathfrak{p}}$. Specifically, an element $\sigma\in C_{k}$ can be written as a finite sum: $\sigma=c_{1}\sigma_{1}+c_{2}\sigma_{2}+\dots+c_{n}\sigma_{n}$, where $\sigma_{i}$ are $k$-simplices, and $c_{i}\in\mathbb{Z}_{\mathfrak{p}}$ are coefficients subject to modular arithmetic, meaning all addition and scalar multiplication are performed modulo $\mathfrak{p}$. A key property of the boundary operator is that the boundary of a boundary is always zero, i.e., $\partial_{k-1}(\partial_{k}(\sigma))=0$ for any $k$-simplex $\sigma$. This ensures that the boundary of a closed structure, such as a cycle, has no further boundary, which is essential for identifying topological features like cycles.
@@ II.1.3 Persistent Homology
Topological features are “born” when they firstly appear in the complex. For example, a new connected component ($H_{0}$) is born when a point first appears as $\epsilon$ increases. Similarly, a cycle ($H_{1}$) is born when a closed cycle (or loop) forms, and a cavity ($H_{2}$) is born when an enclosed space emerges.
@@
In this article, we used the Ripser [75] package to compute persistent homology (PH), and its time complexity for obtaining barcodes from a point cloud with NN points, including the computation of $H_{0}$, $H_{1}$, and $H_{2}$ homology classes, is $O(N^{4})$ in the worst case [75].
@@ II.2.3 Identifying Rearrangements
To characterize the tendency of particle rearrangement, we define the positional change $p_{\text{hop},p}(t)$ of a given particle $p$ within a time window $t-t_{R/2},t+t_{R/2}] [$79, 80]. If $p_{\text{hop},p}(t)$ exceeds a threshold $p_{c}$ during this interval, the particle is labeled as “soft”; otherwise, it is labeled as “hard” [81]. Specifically, $p_{\text{hop},p}(t)$ is defined by:
@@
where $\mathbb{r}_{p}$ is the position vector of particle $p$, and $\langle.\rangle_{A}$ and $\langle.\rangle_{B}$ represent averages over the intervals $A=[t-t_{R/2},t]$ and $B=[t,t+t_{R/2}]$, respectively.
@@ II.3.3 ML Models
[…]
@@
[…]
@@ II.3.4 Scalar Field Trained by Linear SVM
[…]
@@
[…]
@@ II.4 Descriptive Statistics on PH (non-ML approach)
We define a mapping from the results of PH on point cloud data to a real number in the range $0,+\infty)$, referred to as the Separation Index (SI).
@@
Figure 4: a) is the persistence diagram (PD) for a crystal ($\text{Traj}(T^{(1)}_{5})$ at time step $t=900$), of which the homology classes of $H_{0}$ are distributed along a line near the y-axis, while the $H_{1}$ and $H_{2}$ classes are located within the yellow circle and magnified in b). The yellow solid line in b) divides the point sets of $H_{1}$ and $H_{2}$ into two disjoint parts. Inspired by this, the Separation Index (SI) is defined to measure the clarity of the boundary of the point sets for $H_{1}$ and $H_{2}$.
@@ IV.2 Descriptive Statistics on PH (non-ML approach)
In Sec. II.4, we introduced a mapping from the results of the PH analysis to a single real number in the range $0,+\infty)$, specifically the Separation Index (SI), as defined in Eq.9. Since both the neighborhood of a particle and the entire system can be represented as point clouds, the SI applies to both local and global structures, mapping each to a single real number.
@@
Figure 7: The $H_{1}$ and $H_{2}$ points for the entire systems from selected trajectories as follow: a) for liquid ($\text{Traj}(T^{(2)}_{9})$ at time step $t=900$); b) for crystal ($\text{Traj}(T^{(1)}_{5})$ at time step $t=900$); c) for amorphous ($\text{Traj}(T^{(2)}_{1})$ at time step $t=200$).
@@ IV.3 Metrics on Trajectories
Figure 9: The a) Separation Index $\text{SI}(t)$ and b) Global Softness $\mathcal{S}(t)$, where $t=100+10\times(i-1)$, $i\in[1,81]\cap\mathbb{N}$, for trajectories in Group 2, denoted as $\text{Traj}(T^{(2)}_{k}),k\in[1,10]\cap\mathbb{N}$. These 10 trajectories are divided into two categories: those with $k\leq 6$ (Category I) crystallized, while those with $k\geq 7$ (Category II) remained liquid. The metrics over time for each trajectory are shown as faded dashed lines, with bold lines indicating the average in each category—blue for Category I and orange for Category II.
@@
We tracked the evolution in four metrics including normalized MSD, $Q_{6}(t)/\mathscr{N}$, $\text{SI}(t)$ and $\mathcal{S}(t)$ over time on ten trajectories in test group (Group 2), i.e., $\text{Traj}(T^{(2)}_{k}),k\in[1,10]\cap\mathbb{N}$, as well as two trajectories in Group 3, i.e., $\text{Traj}_{\text{KA}}^{(1)}$ and $\text{Traj}_{\text{KA}}^{(2)}$. The data for all metrics are provided in Fig. 19 of Appendix E for reference.
@@ A.2 The Parameter Selection for Rearrangement Tendency (Task 1)
[…]
@@
Figure 16: The histogram of the value of $p_{\text{hop}}$ in $\text{Traj}(T^{(1)}_{6})$ and $\text{Traj}(T^{(1)}_{7})$. The onset temperature $T^{(1)}_{c}$ of crystallization separates $T^{(1)}_{6}$ and $T^{(1)}_{7}$, i.e., $T^{(1)}_{6}<T^{(1)}_{c}<T^{(1)}_{7}$. The red-dotted line is the selected $p_{c}=0.1$. The reason for setting $p_{c}=0.1$ is when $p_{\text{hop}}<p_{c}$, the frequency of $p_{\text{hop}}$ in $\text{Traj}(T^{(1)}_{6})$ is larger than that in $\text{Traj}(T^{(1)}_{7})$; when $p_{\text{hop}}\geq p_{c}$, the frequency of $p_{\text{hop}}$ in $\text{Traj}(T^{(1)}_{6})$ is smaller than that in $\text{Traj}(T^{(1)}_{7})$. $p_{c}=0.1$ is the value at which the relative size of two frequency of $\text{Traj}(T^{(1)}_{6})$ and $\text{Traj}(T^{(1)}_{7})$ reverse.
To select $t_{R/2}$, we use the lowest temperature $T^{(1)}_{1}$ as a reference because it corresponds to the system’s lowest mobility. The interval $t-t_{R/2},t+t_{R/2}]$ should encompass as many complete rearrangements as possible at this low mobility, as shown in Fig. 14. The yellow region marks a full rearrangement, where $p_{\text{hop}}$ rises above and then falls below $p_{c}$. As seen in Fig. 15, most rearrangements at this reference temperature occur within $t_{R}=40$, so we set $t_{R/2}=20$.
As shown in Fig. 16, we chose $p_{c}=0.1$ because it effectively separates the majority of particles in two trajectories above and below the crystallization onset temperature, i.e., $T^{(1)}_{6}$ and $T^{(1)}_{7}$. The $p_{\text{hop},i}(t)$ values are mostly significantly greater or less than $p_{c}=0.1$, respectively. Specifically, $T^{(1)}_{6}$ is the highest temperature at which crystallization occurs, while $T^{(1)}_{7}$ is the lowest temperature where the liquid phase is maintained, both of which are critical for this analysis. By plotting the histograms of $p_{\text{hop}}$ in $\text{Traj}(T^{(1)}_{6})$ and $\text{Traj}(T^{(1)}_{7})$, we determined that $p_{c}=0.1$ is the value at which the relative sizes of the two distributions reverse.
@@ Appendix C Details of Feature Importance Analysis (FIA)
The single decisive predictor mentioned in Sec. IV.1.3, corresponding to the largest eigenvalue from the PCA, is the 1389th element of $\mathbb{I}_{\text{entire}}^{(P)}$, located at pixel (28, 34) in the PI. Given the PI dimensions of $40\times 40$ and birth and persistence ranges of [0.1,1.5][-0.1,1.5], each pixel has a resolution of $0.04$. We can get that the pixel $(28,34)$ is with birth range of $[1.02,1.06]$ and persistence ranges of and $[0.1,0.14]$ in the PD, respectively. The feature importance, based on the eigenvalues of the principal components, revealed that this key predictor corresponds to the 1389th column of the original feature matrix, which is the 1389th element of the vector $\mathbb{I}_{\text{entire}}^{(P)}$.
@@
To begin with, we examined the local bond order parameter $q_{6}$, which characterizes the six-fold symmetry of the local particle environment (defined in Sec. II.2.2), to assess its ability to distinguish particles with different rearrangement tendencies. Specifically, we evaluated the performance of $q_{6}$ in differentiating between the state set of $\{\text{soft},\text{hard}\}$, as reported in Fig. 18a). Here, the states of $\{\text{soft},\text{hard}\}$ are determined by $p_{\text{hop}}$ (see Eq. 4).
Furthermore, we investigated the effectiveness of the Separation Index (SI) in distinguishing the degree of six-fold symmetry preservation in the local particle environment. In particular, we analyzed how SI differentiates between the states $\{\text{ordered},\text{disordered}\}$, with the results reported in Fig. 18b). It is important to note that the states $\{\text{ordered},\text{disordered}\}$ are classified according to the values of $q_{6}$, which reflect the level of local structural order around the particles.